UT Oct 09 2022 - Split night with LBTI




Night Info

  • Sunset 17:55
  • Nautical Twilight Ends 18:48
  • Astro Twilight Ends 19:17
  • Night Center 00:07
  • Handover from LBTI to LBTO 00:03
  • Astro Twilight Begins 04:58
  • Nautical Twilight Begins 05:26
  • Sunrise 06:19
  • Moonrise 17:31
  • Moonset 06:06
  • Moon illumination ~99%

Staff

  • OSA: S.Allanson
  • ISp: Mtn
  • Mtn Manager: R. Hannsen
  • ISA/SO: A.Cardwell, B.Rothberg, R.T. Gatto
  • AO: J.Power

Night Plan

ALTA prediction for Tonight

20221008_see_evol_time_Height_20_20000.png

LBTI Science 1st half

  • Handover at 00:03 - split the reconfig time between LBTI and LUCI-Bino

AO On-Sky Checks

  • Reconfigure for LUCI-Bino
  • Below is not neccesarily order of operations - 1 D-time target is end of night, ESM can be done earlier, etc - depend on conditions in real time
  • OSCO
  • File: Service_Observing/2022B_D_E/LUCI/AO/AO_Testing_V99.xml
    • REQUIRES NEW VERSION OF OT if any sub-frame scripts are generated
  • Thin clouds, seeing 1.2" or better
  • Bin 4
    • Bin 4 test to replicate IT 8735
    • OT script objects
      • RA 23-00hr Star #4
      • RA 02-03hr Star #4
    • NS target Didymos - 1hr
  • Bin 2
    • OT Script Objects
      • RA 23-00hr Star #6
      • RA 02-03hr Star #6

D-Time AO Observations

  • AO/ESM Observations of JWST PSF stars
    • Script under preparation
    • Target 1 RA 03hr Dec -18 - visible 00:15-04:45 - optimal 02:15 Full AO
    • Target 2 RA 03hr Dec -18- visible 00:15-04:45 - optimal 02:15 ESM
    • Target 4 RA 10hr Dec +21 -visible 04:45-twilight (end of night target) Full AO

ESM On-Sky Checks

  • Scripts already extracted - OT is here: Service_Observing/2022B_D_E/LUCI/ESM/ESM_V44.xml
  • M52 - visible twilight-04:15
    • Service_Observing/2022B_D_E/LUCI/ESM/M52
    • Bin1
      • M52_R10pt7_ESM_L1.xml - mono
      • M52_R10pt7_ESM_SL.xml - mixed mode
    • Bin4
      • M52_Bin4_ESM_L1.xml - mono
      • M52_Bin4_ESM_SL.xml - mixed mode
  • Arp 187
    • Service_Observing/2022B_D_E/LUCI/ESM/Arp187
      • Arp187_ESM_L1.xml - mono
      • Arp187_ESM_SL.xml - mixed mode
    • Bin 2 - visible 01:15-twlight
  • NS Target Didymos
    • Bin 4

Poor Weather Backup

  • Photometric Standards (LUCI) - only if clear - any seeing
    • /home/lbto/Service_Observing/2022B_D_E/PoorWeather/LUCI_Phot_ZP_OT_V6.xml
  • G200 Test on-sky
    • /home/lbto/Service_Observing/Restart_2022/LUCI/LUCI_2022_Restart_V1.83.xml
    • any conditions
    • PKS 2247+14 - bright QSO point source - twilight-01:45hr
    • NGC 7027 (PN) - twlight-01:15hr
    • 3C48 - bright QSO point souce - 20hr-twilight
  • ESM - M52 (poor weather checks)


Night Log

Closed Dome

  • UT 01:47 - closed for dark clouds and high humidity - internet lost at UT 23:50 (-1 day) - still out

-- BarryRothberg - 08 Oct 2022
Topic revision: r4 - 09 Oct 2022, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback